OpenStudy (anonymous):

Wite a simplified expression in simplest form to respresent the area of a square

OpenStudy (mokeira):

\[a=s ^{2}\]

OpenStudy (anonymous):

A square has four right angles and all the sides of a square have the same length. Now, if the length of each side is L, the area, A of the square will be: A = L^2.......!!

OpenStudy (anonymous):

area = length * width or a = l *w or a = l * l or a = l^2